OpenStudy (anonymous):

In a garden there are 10 rows and 12 columns of mango trees. The distance between each tree is 2m and distance of 1 m is left from all sides of the boundary of the garden. The length of the garden is

OpenStudy (anonymous):

@ParthKohli need help ...pls help.

OpenStudy (anonymous):

@apoorvk please help out

OpenStudy (anonymous):

Aha, are you preparing for CAT?

OpenStudy (anonymous):

no ... i am in 7th standard

OpenStudy (apoorvk):

so there are 9 gaps of 2m each between the 10 rows. so total length including just the trees --> 9*2 = 18m plus, 1 m boundary on each side = 2m so, total length = 18 + 2 = 20m P.S. - It's not clear what the length actually means - the dimension along the columns or that along the rows. - so check that out.

OpenStudy (anonymous):

The answer is 24m right?

OpenStudy (anonymous):

No bro ans. is 24m @FoolForMath is right pls can you explain

OpenStudy (anonymous):

\[11\times 2+2 =24 \]

OpenStudy (anonymous):

I think apporvk found the breadth.

OpenStudy (anonymous):

How did you get 11?@FoolForMath

OpenStudy (anonymous):

12 columns means 11 gaps.

OpenStudy (anonymous):

You need to find width, so dont focus on rows. There are 10 columns, means 10 trees ( one tree in each column). The distance between two trees is 2 meters. It means from first tree to last tree the total distance is 20 meter ( 10*2) and the left out distance of one meter from both the sides ( 2 meters). Therefore the total distance is 20 + 2 = 22 meters.